# How Does Message Passing Improve Collaborative Filtering?

### v1

@@ 5. Experiments
Amazon-Book 10.30±${}_{\pm{0.21}}$ 12.76±${}_{\pm{0.18}}$ ${}_{\pm{0.18}}$ 11.35±${}_{\pm{0.09}}$ 4.2% 12.67±${}_{\pm{0.06}}$ 12.97±${}_{\pm{0.06}}$ 2.4% 8.01±${}_{\pm{0.25}}$ 8.53±${}_{\pm{0.27}}$ 6.5%
Anime 28.12±${}_{\pm{0.22}}$ 32.82±${}_{\pm{0.21}}$ 34.10±${}_{\pm{0.25}}$ 34.48±${}_{\pm{0.23}}$ 1.1% 29.15±${}_{\pm{0.09}}$ 31.95±${}_{\pm{0.05}}$ 6.9% 35.87±${}_{\pm{0.39}}$ 37.01±${}_{\pm{0.39}}$ 3.2%
Gowalla 17.93±${}_{\pm{0.06}}$ 18.65±${}_{\pm{0.14}}$ 9.68±${}_{\pm{0.06}}$ 9.74±${}_{\pm{0.09}}$ 0.6% 18.30±${}_{\pm{0.17}}$ 18.53±${}_{\pm{0.11}}$ 1.3% 15.93±${}_{\pm{0.21}}$ 16.36±${}_{\pm{0.22}}$ 2.7%
Yelp-2018 10.02±${}_{\pm{0.06}}$ 10.98±${}_{\pm{0.10}}$ 6.89±${}_{\pm{0.09}}$ 7.05±${}_{\pm{0.03}}$ 2.3% 10.81±${}_{\pm{0.10}}$ 11.21±${}_{\pm{0.09}}$ 3.7% 8.41±${}_{\pm{0.19}}$ 9.89±${}_{\pm{0.20}}$ 17.6%
MovieLens-1M 23.93±${}_{\pm{0.14}}$ 26.30±${}_{\pm{0.20}}$ 21.31±${}_{\pm{0.19}}$ 23.88±${}_{\pm{0.25}}$ 12.1% 26.30±${}_{\pm{0.14}}$ 28.40±${}_{\pm{0.15}}$ 8.0% 27.14±${}_{\pm{0.19}}$ 29.78±${}_{\pm{0.23}}$ 9.7%
Internal 6.91±${}_{\pm{0.04}}$ 13.89±${}_{\pm{0.06}}$ OOM - - 11.83±${}_{\pm{0.02}}$ 14.41±${}_{\pm{0.08}}$ 21.8% OOM - -
5.1. Experimental Settings
Baselines. We compare TAG-CF with two branches of methods: (1) CF methods that do not explicitly utilize graphs, including vanilla matrix factorization (MF) methods trained by BPR and DirectAU (rendle2009bpr, ; wang2022towards, ), Efficient Neural Matrix Factorization (chen2020efficient, ) (denoted as ENMF), and UltraGCN (mao2021ultragcn, ). (2) Graph-based CF methods, including LightGCN (he2020lightgcn, ) and NGCF (wang2019neural, ). Besides, we also compare with recent graph-based CF methods that extend LightGCN by adding additional self-supervised signals for better performance, including LightGCL (cai2023lightgcl, ), SimGCL (yu2022graph, ), and SGL (wu2021self, ). For the coherence of reading, we include comprehensive discussions about evaluation protocols across all methods, tuning for hyper-parameters, and other implementation details in Appendix C.
5.2. Performance Improvement to Matrix Factorization Methods
@@
Table 4. The running time and performance of graph-based CF methods that extend LightGCN.
@@
Anime 27.02±${}_{\pm{0.05}}$ 30.48±${}_{\pm{0.12}}$ 28.34±${}_{\pm{0.16}}$ 27.25±${}_{\pm{0.03}}$
Yelp 5.67±${}_{\pm{0.04}}$ 5.99±${}_{\pm{0.09}}$ 4.93±${}_{\pm{0.06}}$ 6.36±${}_{\pm{0.03}}$
Gowalla 9.67±${}_{\pm{0.17}}$ 10.32±${}_{\pm{0.06}}$ 8.99±${}_{\pm{0.13}}$ 9.88±${}_{\pm{0.04}}$
Book 6.69±${}_{\pm{0.02}}$ 7.02±${}_{\pm{0.05}}$ 5.83±${}_{\pm{0.08}}$ 8.13±${}_{\pm{0.03}}$
@@
Anime 31.29±${}_{\pm{0.09}}$ 34.93±${}_{\pm{0.14}}$ 33.64±${}_{\pm{0.22}}$ 31.95±${}_{\pm{0.05}}$
Yelp 10.01±${}_{\pm{0.08}}$ 10.56±${}_{\pm{0.13}}$ 8.83±${}_{\pm{0.04}}$ 11.21±${}_{\pm{0.09}}$
Gowalla 18.18±${}_{\pm{0.24}}$ 19.22±${}_{\pm{0.09}}$ 16.99±${}_{\pm{0.10}}$ 18.53±${}_{\pm{0.09}}$
Book 11.15±${}_{\pm{0.04}}$ 11.51±${}_{\pm{0.09}}$ 10.06±${}_{\pm{0.05}}$ 12.97±${}_{\pm{0.06}}$
@@
Table 5. Performance of TAG-CF when applied to models trained with BPR loss.
@@
Anime 30.02±${}_{\pm{0.07}}$ 29.36±${}_{\pm{0.23}}$ 30.56±${}_{\pm{0.27}}$ 4.1%
Yelp 4.34±${}_{\pm{0.07}}$ 3.63±${}_{\pm{0.15}}$ 3.81±${}_{\pm{0.18}}$ 5.0%
Gowalla 8.22±${}_{\pm{0.03}}$ 7.56±${}_{\pm{0.14}}$ 7.88±${}_{\pm{0.15}}$ 4.2%
Book 5.19±${}_{\pm{0.14}}$ 4.19±${}_{\pm{0.14}}$ 4.68±${}_{\pm{0.14}}$ 11.7%
@@
Anime 30.14±${}_{\pm{0.07}}$ 29.51±${}_{\pm{0.21}}$ 30.23±${}_{\pm{0.26}}$ 2.4%
Yelp 4.87±${}_{\pm{0.06}}$ 3.96±${}_{\pm{0.14}}$ 4.26±${}_{\pm{0.17}}$ 7.6%
Gowalla 8.32±${}_{\pm{0.03}}$ 7.51±${}_{\pm{0.12}}$ 7.99±${}_{\pm{0.14}}$ 6.4%
Book 5.07±${}_{\pm{0.15}}$ 4.15±${}_{\pm{0.13}}$ 4.32±${}_{\pm{0.13}}$ 4.1%
@@
Anime 34.23±${}_{\pm{0.08}}$ 34.81±${}_{\pm{0.32}}$ 35.42±${}_{\pm{0.35}}$ 1.8%
Yelp 8.19±${}_{\pm{0.20}}$ 6.93±${}_{\pm{0.26}}$ 7.25±${}_{\pm{0.19}}$ 4.6%
Gowalla 16.17±${}_{\pm{0.12}}$ 14.86±${}_{\pm{0.23}}$ 15.33±${}_{\pm{0.24}}$ 3.2%
Book 8.81±${}_{\pm{0.26}}$ 7.45±${}_{\pm{0.22}}$ 8.05±${}_{\pm{0.15}}$ 8.1%
@@
Anime 34.21±${}_{\pm{0.08}}$ 34.84±${}_{\pm{0.30}}$ 35.23±${}_{\pm{0.34}}$ 1.1%
Yelp 8.33±${}_{\pm{0.30}}$ 7.27±${}_{\pm{0.27}}$ 7.62±${}_{\pm{0.22}}$ 4.8%
Gowalla 15.69±${}_{\pm{0.07}}$ 14.47±${}_{\pm{0.23}}$ 14.92±${}_{\pm{0.25}}$ 3.1%
Book 8.65±${}_{\pm{0.24}}$ 7.35±${}_{\pm{0.22}}$ 7.64±${}_{\pm{0.20}}$ 3.9%
5.3. Performance Comparison Among Graph-based Methods
@@
We further compare TAG-CF with state-of-the-art graph-based CF methods, with their performance and efficiency shown in Table 4. Among these performant baselines, TAG-CF exhibits competitive performance, with an average rank of 1.2 on NDCG and 1.7 on Recall. Though not always the model that delivers the best performance, TAG-CF can deliver comparably promising results and introduces little computational overheads (i.e., ranked 1.0 for running time). Considering efficiency as one factor, TAG-CF achieves the best performance across all baselines with an average rank of 1.9.
@@
5.4. Effectiveness for Different Training Signals
To answer RQ (3), besides DirectAU, we also conduct experiments on BPR loss, as shown in Table 5. When applied to BPR, TAG-CF still consistently improves the performance by large margins (i.e., 6.3% and 5.1% average improvement on low-degree and overall NDCG respectively, and 4.4% and 3.2% on low-degree and overall Recall respectively). We notice that TAG-CF sometimes does not perform as competitively as LightGCN when both are trained with BPR. We check norms of learned representations from MF with BPR and discover that they have high variance since BPR does not explicitly enforce any regularization. This might not favor TAG-CF as a test-time augmentation method due to its simple design, which cannot adapt representations with high variance.
5.5. Performance w.r.t. User Degree
Figure 2. The performance and efficiency improvement of ${}^{+}$ w.r.t. different cutoffs. ${}^{+}$ further improves TAG-CF with less computational overheads. 100% is the original performance/efficiency of vanilla TAG-CF.
To answer RQ (4), we apply ${}^{+}$ to four public datasets and the performance and the efficiency improvement are demonstrated in Figure 2. Overall, the running time improvement brought by ${}^{+}$ exponentially increases as the degree decreases, since low-degree users have sparse neighborhoods and there is hence less information for ${}^{+}$ to aggregation. When the degree cutoff is low (i.e., less than 100), the effectiveness of ${}^{+}$ proportional increases as the degree cutoff increases.
When setting the cutoff to a user degree of around 100, on Amazon-Book, Gowalla, and Yelp-2018, ${}^{+}$ can further improve TAG-CF by 125%, 17%, and 11%, respectively, with efficiency improvement of 7%, 4%, and 8%. In these cases, ${}^{+}$ not only significantly improves the performance but also effectively reduces computational overheads. However, on these three datasets, after the cutoff bypasses a degree of 100, the performance improvement eventually decreases to the performance of TAG-CF (i.e., 100%), indicating that test-time aggregation jeopardizes the performance on high-degree nodes. On Anime, though no downgrade on high-degree users, the performance improvement of ${}^{+}$ to TAG-CF is incremental. These phenomenons not only demonstrate the effectiveness and efficiency of ${}^{+}$, but also verify our findings in Section 3.2 that message passing in CF helps low-degree users more than high-degree users.
Table 6. Improvement (NDCG@20) brought by TAG-CF at different degree cutoffs and upsampling rates on ML-1M.
Up-sampling Up-sample Rate: 100% Up-sample Rate: 300%
Degree
MF + TAG-CF Impr. ($\uparrow$%) MF + TAG-CF Impr. ($\uparrow$%)
40 20.62 28.87 38.8% 19.30 25.01 30.3%
80 20.10 27.43 35.9% 18.40 23.30 26.8%
160 19.39 26.63 36.6% 17.93 23.37 29.8%
5.6. Analysis of TAG-CF through Up-sampling
6. Conclusion
@@ Appendix A Proof of Theorem 1
One preliminary theoretical foundation for Theorem 1 to hold is that a K-layer graph convolution network (GCN) exactly optimizes the second term in Equation 8, which has been proved by (zhu2021interpreting, ). With this preliminary, the proof to Theorem 1 starts as:
@@
(10) $\mathcal{L}_{\text{DirectAU}}$ $=\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$
(11) $+\sum_{u,u^{\prime}\in\mathcal{U}}\log e^{-2||\mathbf{u}-\mathbf{u}^{\prime}||^{2}}+\sum_{i,i^{\prime}\in\mathcal{I}}\log e^{-2||\mathbf{i}-\mathbf{i}^{\prime}||^{2}}.$
@@
(12) $\mathcal{L}_{\text{BPR}}$ $=-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(s_{ij}-s_{ik})=$
(13) $-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k})$
(14) $=\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}-\log\Big{(}\frac{e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}}{e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}}\Big{)}$
(15) $=\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}+\log\Big{(}e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}\Big{)}$
@@
(16) $\mathcal{L}_{\text{BPR}}=\frac{1}{2}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}-1+\log\Big{(}e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}\Big{)}.$

### v2

@@ 5.2 Performance Improvement to Matrix Factorization Methods
Amazon-Book 10.30±0.21 12.76±0.18 10.89±0.18 11.35±0.09 4.2% 12.67±0.06 12.97±0.06 2.4% 8.01±0.25 8.53±0.27 6.5%
Anime 28.12±0.22 32.82±0.21 34.10±0.25 34.48±0.23 1.1% 29.15±0.09 31.95±0.05 6.9% 35.87±0.39 37.01±0.39 3.2%
Gowalla 17.93±0.06 18.65±0.14 9.68±0.06 9.74±0.09 0.6% 18.30±0.17 18.53±0.11 1.3% 15.93±0.21 16.36±0.22 2.7%
Yelp-2018 10.02±0.06 10.98±0.10 6.89±0.09 7.05±0.03 2.3% 10.81±0.10 11.21±0.09 3.7% 8.41±0.19 9.89±0.20 17.6%
MovieLens-1M 23.93±0.14 26.30±0.20 21.31±0.19 23.88±0.25 12.1% 26.30±0.14 28.40±0.15 8.0% 27.14±0.19 29.78±0.23 9.7%
Internal 6.91±0.04 13.89±0.06 OOM - - 11.83±0.02 14.41±0.08 21.8% OOM - -
@@
Table 4: The running time and performance of graph-based CF methods that extend LightGCN.
@@
Anime 27.02±0.05 30.48±0.12 28.34±0.16 27.25±0.03
Yelp 5.67±0.04 5.99±0.09 4.93±0.06 6.36±0.03
Gowalla 9.67±0.17 10.32±0.06 8.99±0.13 9.88±0.04
Book 6.69±0.02 7.02±0.05 5.83±0.08 8.13±0.03
@@
Anime 31.29±0.09 34.93±0.14 33.64±0.22 31.95±0.05
Yelp 10.01±0.08 10.56±0.13 8.83±0.04 11.21±0.09
Gowalla 18.18±0.24 19.22±0.09 16.99±0.10 18.53±0.09
Book 11.15±0.04 11.51±0.09 10.06±0.05 12.97±0.06
@@
Table 5: Performance of TAG-CF when applied to models trained with BPR loss.
@@
Anime 30.02±0.07 29.36±0.23 30.56±0.27 4.1%
Yelp 4.34±0.07 3.63±0.15 3.81±0.18 5.0%
Gowalla 8.22±0.03 7.56±0.14 7.88±0.15 4.2%
Book 5.19±0.14 4.19±0.14 4.68±0.14 11.7%
@@
Anime 30.14±0.07 29.51±0.21 30.23±0.26 2.4%
Yelp 4.87±0.06 3.96±0.14 4.26±0.17 7.6%
Gowalla 8.32±0.03 7.51±0.12 7.99±0.14 6.4%
Book 5.07±0.15 4.15±0.13 4.32±0.13 4.1%
@@
Anime 34.23±0.08 34.81±0.32 35.42±0.35 1.8%
Yelp 8.19±0.20 6.93±0.26 7.25±0.19 4.6%
Gowalla 16.17±0.12 14.86±0.23 15.33±0.24 3.2%
Book 8.81±0.26 7.45±0.22 8.05±0.15 8.1%
@@
Anime 34.21±0.08 34.84±0.30 35.23±0.34 1.1%
Yelp 8.33±0.30 7.27±0.27 7.62±0.22 4.8%
Gowalla 15.69±0.07 14.47±0.23 14.92±0.25 3.1%
Book 8.65±0.24 7.35±0.22 7.64±0.20 3.9%
5.3 Performance Comparison Among Graph-based Methods
@@
We further compare TAG-CF with state-of-the-art graph-based CF methods, with their performance and efficiency shown in Section 5.2. Among these performant baselines, TAG-CF exhibits competitive performance, with an average rank of 1.2 on NDCG and 1.7 on Recall. Though not always the model that delivers the best performance, TAG-CF can deliver comparably promising results and introduces little computational overheads (i.e., ranked 1.0 for running time). Considering efficiency as one factor, TAG-CF achieves the best performance across all baselines with an average rank of 1.9.
@@
5.4 Effectiveness for Different Training Signals
To answer RQ (3), besides DirectAU, we also conduct experiments on BPR loss, as shown in Section 5.2. When applied to BPR, TAG-CF still consistently improves the performance by large margins (i.e., 6.3% and 5.1% average improvement on low-degree and overall NDCG respectively, and 4.4% and 3.2% on low-degree and overall Recall respectively). We notice that TAG-CF sometimes does not perform as competitively as LightGCN when both are trained with BPR. We check norms of learned representations from MF with BPR and discover that they have high variance since BPR does not explicitly enforce any regularization. This might not favor TAG-CF as a test-time augmentation method due to its simple design, which cannot adapt representations with high variance.
5.5 Performance w.r.t. User Degree
To answer RQ (4), we apply TAG-CF+ to four public datasets and the performance and the efficiency improvement are demonstrated in Figure 2. Overall, the running time improvement brought by TAG-CF+ exponentially increases as the degree decreases, since low-degree users have sparse neighborhoods and there is hence less information for TAG-CF+ to aggregation. When the degree cutoff is low (i.e., less than 100), the effectiveness of TAG-CF+ proportional increases as the degree cutoff increases.
When setting the cutoff to a user degree of around 100, on Amazon-Book, Gowalla, and Yelp-2018, TAG-CF+ can further improve TAG-CF by 125%, 17%, and 11%, respectively, with efficiency improvement of 7%, 4%, and 8%. In these cases, TAG-CF+ not only significantly improves the performance but also effectively reduces computational overheads.
Figure 2: The performance and efficiency improvement of TAG-CF+ w.r.t. different cutoffs. TAG-CF+ further improves TAG-CF with less computational overheads. 100% is the original performance/efficiency of vanilla TAG-CF.
However, on these three datasets, after the cutoff bypasses a degree of 100, the performance improvement eventually decreases to the performance of TAG-CF (i.e., 100%), indicating that test-time aggregation jeopardizes the performance on high-degree nodes. On Anime, though no downgrade on high-degree users, the performance improvement of TAG-CF+ to TAG-CF is incremental. These phenomenons not only demonstrate the effectiveness and efficiency of TAG-CF+, but also verify our findings in Section 3.2 that message passing in CF helps low-degree users more than high-degree users.
6 Conclusion
@@
7 Limitation and Broader Impact
One limitation of our proposal could be the utilization of graphs in large-scale machine learning pipeline. TAG-CF conducts a single-time aggregation of neighbors, which could be equivalently achieved by existing technologies such as SQL, BigQuery, etc. Furthermore, we observe no ethical concern entailed by our proposal, but we note that both ethical or unethical applications based on collaborative filtering may benefit from the effectiveness of our work. Care should be taken to ensure socially positive and beneficial results of machine learning algorithms.
8 Acknowledgments
This work was mostly conducted during the internship of Clark, William, and Zhichun at Snap Inc. We would like to thank Xin Chen and his colleagues from Snap Inc. for their help on pre-processing the internal dataset. This work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2203262, IIS-2217239, CNS-2426514, CNS-2203261, and CMMI-2146076. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
We sincerely appreciate constructive feedback from all reviewers during the paper review phase.
@@ Appendix A Proof of Theorem 1
Here we re-state Theorem 1 before diving into its proof:
Theorem 1. Given that a K-layer GCN minimizes $\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$, During the training of MF methods, assuming that $||\mathbf{u}_{i}||^{2}=||\mathbf{i}_{j}||^{2}=1$ for any $u_{i}\in\mathcal{U}$ and $I_{j}\in\mathcal{I}$, objectives of BPR and DirectAU are strictly upper-bounded by the objective of message passing (i.e., $\mathcal{L}_{\text{BPR}}\leq\sum_{(i,j)\in\mathcal{D}}$ $||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$ and $\mathcal{L}_{\text{DirectAU}}\leq\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$).
One preliminary theoretical foundation for Theorem 1 to hold is that a K-layer graph convolution network (GCN) exactly optimizes the second term in Equation 6, which has been proved by [82]. For ease of reading, we re-phrase it again as the following:
Theorem 2.
The message passing for GCN optimizes the following graph regularization term: $\mathcal{O}=\min_{\mathbf{Z}}\{tr(\mathbf{Z}^{\intercal}\mathbf{L}\mathbf{Z}))\}$.
Proof.
Set derivative of $tr(\mathbf{Z}^{\intercal}\mathbf{L}\mathbf{Z})$ with respect to $\mathbf{Z}$ to zero:
[…]
$\mathbf{Z}^{(K)}=\mathbf{A}\mathbf{Z}^{(K-1)}$ (9)
which indicates:
$\mathbf{Z}^{(K)}=\mathbf{A}\mathbf{Z}^{(K-1)}=\mathbf{A}^{2}\mathbf{Z}^{(K-2)}=\dots=\mathbf{A}^{K}\mathbf{Z}^{(0)}=\mathbf{A}^{K}\mathbf{X}\mathbf{W}.$ (10)
According to this theoretical foundation, it is straightforward that Theorem 2 is also applicable for the message passing of LightGCN in the setting of CF if we let $\mathbf{A}=\{0,1\}^{(|\mathcal{U}|+|\mathcal{I}|)\times(|\mathcal{U}|+|\mathcal{I}|)}$, $\mathbf{X}=\mathbf{I}_{|\mathcal{U}|+|\mathcal{I}|}$, and $\mathbf{W}=(\mathbf{U}||\mathbf{I}$), where ||||| | refers to the concatenation operation. With this preliminary, the proof to Theorem 1 starts as:
@@
$\mathcal{L}_{\text{DirectAU}}$ $=\sum_{(i,j)\in\mathcal{D}}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}$ (11)
$+\sum_{u,u^{\prime}\in\mathcal{U}}\log e^{-2||\mathbf{u}-\mathbf{u}^{\prime}||^{2}}+\sum_{i,i^{\prime}\in\mathcal{I}}\log e^{-2||\mathbf{i}-\mathbf{i}^{\prime}||^{2}}.$ (12)
@@
$\mathcal{L}_{\text{BPR}}$ $=-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(s_{ij}-s_{ik})=$ (13)
$-\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}\log\sigma(\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k})$ (14)
$=\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}-\log\Big{(}\frac{e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}}{e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}}\Big{)}$ (15)
$=\sum_{(i,j)\in\mathcal{D}}\sum_{(i,k)\notin\mathcal{D}}-\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}+\log\Big{(}e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}\Big{)}$ (16)
@@
$\mathcal{L}_{\text{BPR}}=\frac{1}{2}||\mathbf{u}_{i}-\mathbf{i}_{j}||^{2}-1+\log\Big{(}e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{j}}+e^{\mathbf{u}_{i}^{\intercal}\cdot\mathbf{i}_{k}}\Big{)}.$ (17)